# Perry: A High-level Framework for Accelerating Cyber Deception Experimentation

### v1

@@ IV-C Observability module
Similar to actions, defenders and attackers need to reason about the state of an environment (e.g., host used a decoy credential, infected a new host, found critical data). As seen in Sec. II, existing tools expose low-level events (e.g., system calls, network tracing, outputs of a bash command, etc.) and require environment specific information (e.g., IP address of a decoy host). To make it easier to write and modify high-level strategies, we design an environment agnostic observability module inspired by Zeek [46] that decouples strategies’ high-level observations from their low-level events.
Current attack and defense emulation tools require implementations to have parsers for low-level telemetry data (e.g., network traces, command outputs) [5, 23, 15, 14, 25]. As a result, strategy implementations become (1) hard to modify because they have large amounts of low-level code to process the data, and (2) hard to extend because they are tightly coupled to the type of telemetry data.
@@
Here, we decouple the high-level intention of detecting hosts using decoy credentials from the low-level implementation. Consequently, we can quickly explore new telemetry techniques without modifying the strategy. For instance, if we implement a new anomaly detection algorithm to output HostUsedDecoyCredential events, it works with the existing strategy. We design a library of common high-level deception and attacker events shown in Table II.
@@ IV-D Attack graph service
Attack and defense strategies often have to reason about vulnerabilities in networks. For example, an attack may want to target a specific host or get privileged access to an already infected host. Defenders too may want to proactively explore possible future attack steps. As seen in Sec. II, existing tools only offer low-level methods for reasoning about potential targets in a network. To make it easier to reason about vulnerabilities, we design an attack graph service that decouples strategies and from their complex reasoning of vulnerabilities.
@@
Because attackers often have partial knowledge about the environment, the attack graph service continually updates itself as new information is discovered using the environment state service. For example, if credentials were discovered on a host, the attack graph service will identify new paths that infect hosts. These paths can directly be executed by the action translator. As a result, the attack graph service removes complex logic from the strategy and creates a modular query so that other strategies can also request infection methods.
@@ V Illustrative example
In this section, we show how using Perry can simplify the implementation for the stateful deception approach and Equifax attacker in Fig. 5. Then, we illustrate how operators can extend Perry through several qualitative examples.
@@
In the initialization phase, we use the environment state service to randomly select subnets to deploy decoy hosts to and randomly select hosts to deploy decoy credentials to. The action translator then translates the deploy decoy and credential actions into their low-level commands. We use the observability module to subscribe to a key event, a decoy host interaction. Next, we specify the event handlers to react to these interactions by restoring both the decoy host and the host that interacted with the decoy host.
@@
Overall, we are able to reduce the total lines of code from 467 to 44 and from 1596 to 94 for the defender and attacker respectively (Sec. VII-A). Even more importantly, Perry gives an intuitive and high-level abstraction for the user. In the next section, we discuss the detailed design of the Perry modules.
@@ VI Implementation
[…]
The environment’s network topology is specified using Terraform Language [21], an infrastructure as code language. The assets, host configurations, and vulnerabilities, are specified using an internal Python SDK. The internal Python SDK is designed to be extensible for users to add additional assets, configurations, and vulnerabilities.
Environments are instantiated in two phases: setup and launch. The setup phase deploys and configures the network and hosts in OpenStack. Next, assets, host configurations, and vulnerabilities are configured through a mixture of Python and Ansible [50]. Once the environment is configured, Perry snapshots each host to save time launching the environment. During the launch phase, Perry restores the network topology to its original state. Then, Perry instantiates each host on the network using the saved snapshots.
@@
The defender’s action translator uses orchestrators to enable Perry to work in multiple environments such as OpenStack [43], or AWS [4]. In this paper, we implement an OpenStack orchestrator. For a high-level action, the orchestrator will use the OpenStack SDK to execute the low-level actions. We use Ansible to execute the remaining low-level actions that are not specific to the cloud environment.
We implement the defender’s observability module with Python and Elasticsearch [14]. The observability module continually continually queries the database and has rules that raise high-level observations.
Attackers: We implement Perry’s attacker abstraction by translating Perry’s high-level attack strategies into low-level Caldera [5] code. The action translator maps high-level actions to low-level Caldera actions. Caldera low-level actions are template files that contain bash commands or other binaries to execute. We implement the observability module by considering the outputs of each low-level action as low-level events. For each low-level action, we create rules that parse these low-level outputs and raise high-level observations. Similar to the defender, we implement a library of common high-level attack actions and events shown in Table I and Table II.
@@ VII Evaluation
In this section, we first evaluate how Perry’s abstractions reduce the implementation effort for a wide variety of defenders and attackers. Second, we discuss how Perry is extensible to enable operators to specify variations of existing strategies and explore new defense and attack capabilities as they become available. Third, we show how Perry can shed light on the efficacy of various defenses and their cost-benefit tradeoffs.
@@ VII-A Reduction in implementation effort
For attack implementations, we compare Perry against implementations using low-level Caldera actions. In Fig. 7, we show Perry reduces LOC of the attacker implementations 16.9–26.9$\times$. The action translator provided the biggest benefit for LOC reduction by offloading the complex logic to execute low-level calls to the Caldera SDK. The Darkside strategy had the largest savings in LOC, with a 26.9$\times$ reduction.
Figure 7: Across all defender strategies, Perry’s abstractions reduce the LOC by 6.0-12.7$\times$. Across all attacker strategies, Perry’s abstractions reduce the LOC by 13.1-22.2$\times$. Most of the reduction is from the action translator.
For defense strategies, we compare Perry vs.  implementations expressed using a combination of Python code and Elasticsearch. In Fig. 7, we show Perry reduces LOC of the deception implementations by 6.8–14.3$\times$. Again, the action translator provides the largest benefit by avoiding the complex interactions with the OpenStack and Ansible SDKs, shown in Fig. 7. The second highest benefit stems from the environment state service, as deception strategies often need to reason about the current state of the environment.
@@
LLM-based code generation: Given recent advances in LLM-based code generation, a natural question is, if and how, the aforementioned effort can be reduced using LLMs. As a first-step, we evaluate LLMs ability at generating three deception defenses. To this end, we consider two settings: (1) prompting an LLM to produce code using current tools and low level APIs and commands, and (2) providing the LLM with Perry APIs and then prompting it to produce code using this API (Fig. 8). For both settings, we use Sonnet 3.7 Thinking as the LLM. The prompts we use are in Appendix  F.
We execute the code to detect errors and also manually review it to validate the implementation. For all three deception strategies, LLMs are able to correctly translate the strategy with Perry’s abstractions. However, the code produced in low-level APIs is incorrect and do not execute without errors. We validated the LLM-generated Perry code against the DFS attacker in the Equifax Large environment and find that the behavior is comparable to the manually written defense code.
@@ VII-B Evaluating Perry’s extensibility
Defender capabilities: As an illustrative case study, we consider the case where an operator wants to explore the benefits of fine-grained host telemetry tools [12, 26] to identify when attackers interact with decoy resources. Specifically, we consider SysFlow [26] to collect system calls from eBPF [12]. This required us to 1) create an Ansible playbook to install SysFlow on each host (26 LOC); 2) add a Python module to configure SysFlow to send the data to the Elasticsearch database (28 LOC); 3) modify environment files to install SysFlow on all hosts (6 LOC); and 4) add an observation module that uses the system calls to detect when attackers use decoy credentials or interact with decoy hosts (42 LOC).
@@
New attacker capabilities: Attackers expressed using the Perry APIs can be extended similar to the defender. Furthermore, Perry can also to support attacker code that does not use the Perry API. For instance, semi- or autonomous LLM-based attackers are starting to show early promise in CTF style challenges [59]. It is easy to plug to plug in such emerging attacker capabilities as well. The effort to support a new autonomous attacker in Perry was quite minimal; we just need to create a new attack strategy that prompts the LLM for bash commands to execute and respond with the results (52 LOC).
@@ VII-C What-If scenarios and interesting findings
Illustrative findings: In what follows, we highlight some interesting findings. We do not intend these as conclusive statements about the state of deception or universal facts across all future environments. We use these as illustrative examples of the types of what-if analysis tradeoffs for operators using Perry. For each result, we run the experiment 5 times and report the averages across the runs. In the interest of brevity, for some results we only focus on the environments inspired by the Equifax and Colonial Pipeline incidents.
@@
In the Equifax-inspired environment  we consider: (1) binary success metric if the attacker was able to exfiltrate all data and (b) the time to exfiltrate all data. In Colonial Pipeline-inspired environment, similarly we use a binary metric if the attacker was able to infect all critical actuators (Fig. 11) and the time the attacker took to infect all actuators (Fig. 9). [^3] For both time metrics, if the attacker failed, we mark the time as the maximum for the environment.
@@
We find static deception strategies can delay but not fundamentally limit the “binary” success (i.e., infect or exfiltrate critical assets) of many attackers we evaluated. In contrast, stateful deception that reacts to changes in system state are able thwart many types of attacks we considered (Fig. 11).
@@
Figure 12: In the Colonial Pipeline-inspired environment against a targeted attacker, a misleading OT placement strategy can slow the attacker down by 3.2$\times$ more than a layered strategy.
@@
We also discover altering the placement policy can further amplify the efficacy slow down the targeted attacker 3.2$\times$. To identify why this placement policy is so effective, we show two example timelines of attacks in Fig. 12(b). The placement policy “floods the zone” of the OT network at the tradeoff of allowing the attacker to gain access to non-critical hosts.
@@
Table VII summarizes running our Sonnet 3.7 Thinking LLM-based attacker for various deception scenarios across 10 trials. (Note that this is different from the use of LLM based code generation for defense strategies in Sec. VII-A). The LLM-based attacker successfully exfiltrated data in 2 out of 10 trials with no deception as shown in Table VII. However, against both deception strategies, it failed to exfiltrate data in any trial. We found that $\approx 90\%$ of commands from the LLM were indeed “wasted” on deceptive resources. While these are the early days in autonomous LLM-based attackers, this early result suggests deception may be a promising defense against future autonomous LLM-based attackers.
@@ VIII Discussion and Limitations
Perry could also potentially be used to generate and collect realistic synthetic data about network attacks to train anomaly detection tools. For example, Perry could run a wide variety of attackers in an environment to train intrusion-detection tools.
@@
Furthermore, a challenge with all emulation tools is the high resource requirements of emulating large networks. For instance, we estimate that emulating a 10,000-host enterprise organization for the same number of experiments in Sec. VII-A would cost \$50,594 on Google Cloud  [^4] We assume each virtual machine has the same requirements as the experiments in Sec. VII-A: 1 CPU, 1GB RAM, and 5GB of disk space. We assume each experiment takes on average 1 hour.. To help reduce the cost of evaluating deception on large networks, we have preliminarily evaluated whether experimental results obtained on smaller, simplified environments can carry over to larger, more realistic environments; our experiments suggest that results obtained on smaller environments can carry over to larger environments. As future work, we plan to more thoroughly investigate how to reliably approximate aspects of the environment such as services, vulnerabilities, and human interactions.
@@ X Conclusions
Our work was inspired by the significant pain points we faced in emulating attackers, deception approaches and environments with existing tools. Our experiences and challenges helped us identify common design patterns to inform Perry’s programming model for implementing deception what-if experiments. We use Perry to demonstrate the importance of what-if experiments in the context of cyber deception. We hope this work will serve as the basis to provide operators with the tools and techniques to explore and optimize deception strategies in their deployment.

### v2

@@ IV-C Observability module
Similar to actions, defenders and attackers need to reason about the state of an environment (e.g., host used a decoy credential, infected a new host, found critical data). As seen in Sec. II, existing tools expose low-level events (e.g., system calls, network tracing, outputs of a bash command, etc.) and require environment-specific information (e.g., IP address of a decoy host). To make it easier to write and modify high-level strategies, we design an environment-agnostic observability module inspired by Zeek [49] that decouples strategies’ high-level observations from their low-level events.
Current attack and defense emulation tools require implementations to have parsers for low-level telemetry data (e.g., network traces, command outputs) [5, 25, 16, 15, 27]. As a result, strategy implementations become (1) hard to modify because they have large amounts of low-level code to process the data and (2) hard to extend because they are tightly coupled to the type of telemetry data.
@@
Here, we decouple the high-level intention of detecting hosts using decoy credentials from the low-level implementation. Consequently, we can quickly explore new telemetry techniques without modifying the strategy. For instance, if we implement a new anomaly detection algorithm to output HostUsedDecoyCredential events, it works with the existing strategy. We design a library of common high-level deception and attacker events, shown in Table II.
@@ IV-D Attack graph service
Attack and defense strategies often have to reason about vulnerabilities in networks. For example, an attack may want to target a specific host or get privileged access to an already infected host. Defenders too may want to proactively explore possible future attack steps. As seen in Sec. II, existing tools only offer low-level methods for reasoning about potential targets in a network. To make it easier to reason about vulnerabilities, we design an attack graph service.
@@
Because attackers can only have partial knowledge about the environment, the attack graph service continually updates itself as new information is discovered using the environment state service. For example, if credentials were discovered on a host, the attack graph service will identify new paths that infect hosts. These paths can directly be executed by the action translator. As a result, the attack graph service removes complex logic from the strategy and creates a modular query so that other strategies can also request infection methods.
IV-E Attacker service
Our attacker service enables us to model realistic attackers. Attackers in Perry use C&C servers, install malware agents on hosts, scan networks, and execute real exploits. The high-level attacker actions in Table I use the action translation module to translate these high-level actions into common low-level attack tools. For instance, the lateral movement high-level action either uses credentials or exploits to install malware agents on the target host. The exploits are currently selected from a small built-in library; however, Perry also supports an optional module that can use Metasploit’s larger exploit library. In addition, we envision the community creating custom implementations of these actions, such as a lateral movement implementation that uses LLMs to generate the exploits [11, 57].
@@ V Illustrative example
In this section, we show how using Perry can simplify the implementation for the stateful deception approach and Equifax attacker in Fig. 5. These examples illustrate the expressiveness of how deception strategies, attackers, and environments are specified in Perry. Then, we illustrate how operators can extend Perry through qualitative examples.
@@
In the initialization phase, we use the environment state service to select random subnets to deploy decoy hosts to and select random hosts to deploy decoy credentials to. The action translator then translates the deploy decoy and credential actions into their low-level commands. We use the observability module to subscribe to a key event, a decoy host interaction. Next, we specify the event handlers to react to these interactions by restoring both the decoy host and the host that interacted with the decoy host.
@@
Overall, we are able to reduce the total lines of code from 467 to 44 and from 1596 to 94 for the defender and attacker respectively (Sec. VII-A). Even more importantly, Perry gives an intuitive high-level abstraction for the user. In the next section, we discuss the detailed design of the Perry modules.
Environment expressiveness: In Perry, environments are also specified at a high level. The network topologies are defined with Terraform, an industry standard for specifying virtual network topologies [23]. In addition to the topology, environments have a Python specification with the environments vulnerabilities, host configurations, and key assets. Below is a snippet of the specification of the Equifax-inspired environment used in Sec. VII:
# (A) Setup ApacheStruts Vulnerability
for web_server in self.webservers:
SetupStrutsVulnerability(web_server, port=443)
# (B) Setup credentials
cred_web_server = self.webservers[0]
for db in self.database_hosts:
SetupServerSSHKeys(
cred_web_server.ip,
cred_web_server.users[0],
db.ip,
db.users[0]
# (C) Setup database data
for db in self.database_hosts:
AddData(db, database.users[0])
In this example, first the vulnerable ApacheStruts web service is installed on the two web servers. Then, one web sever is given SSH credentials to all of the databases. Finally, each data is added to each database for the defender to protect.
@@ VI Implementation
In Perry, we define an environment as a tuple $\langle\mathit{Topology},\mathit{Assets},\textit{Host configuration},\textit{Vulnerabilities}\rangle$. The topology is the routers, switches, firewalls, and hosts of the network. The assets are objects such as critical data. The host configuration are the configurations of hosts such as deploying a database or creating a user. The vulnerabilities misconfigure hosts or setup vulnerable software versions.
The environment’s network topology is specified using Terraform Language [23], an “infrastructure as code” language. The assets, host configurations, and vulnerabilities, are specified using an internal Python SDK. The internal Python SDK is designed to be extensible for users to add additional assets, configurations, and vulnerabilities.
Environments are instantiated in two phases: setup and launch. The setup phase deploys and configures the network and hosts in OpenStack. Next, assets, host configurations, and vulnerabilities are configured through a mixture of Python and Ansible [53]. Once the environment is configured, Perry snapshots each host to save time launching the environment. During the launch phase, Perry restores the network topology to its original state. Then, Perry instantiates each host on the network using the saved snapshots. The VM snapshots are RAW images which are portable to most cloud providers (e.g., AWS, Google Cloud). [^3] As future work, we plan to support several major cloud providers in the open-source repository.
@@
The defender’s action translator uses orchestrators to enable Perry to work in multiple clouds such as OpenStack [46], or AWS [4]. In this paper, we implement an OpenStack orchestrator. For a high-level action, the orchestrator will use the OpenStack SDK to execute the low-level actions. We use Ansible to execute the remaining low-level actions that are not specific to the cloud environment. We implement the defender’s observability module with Python and Elasticsearch [15]. The observability module continually queries the database and has rules that raise high-level observations.
Attackers: The attacker service uses Caldera, an open-source C&C server to communicate and send commands to infected hosts. [^4] Other C&C server’s such as Mythic [44] or CobaltStrike [18] could also be used. The action translator translates tasks from Perry’s high-level attack strategies into low-level commands (e.g., Shell code, Python scripts) that are then executed by Caldera [5]. We implement the observability module by considering the outputs of each low-level action as low-level events. For each low-level action, we create rules that parse these low-level outputs and raise high-level observations. Similar to the defender, we implement a library of common high-level attack actions and events shown in Table I and Table II.
@@ VII Evaluation
In this section, we first evaluate how Perry’s abstractions reduce the implementation effort for a wide variety of deceptive defenders and attackers. Second, we discuss how Perry is extensible to enable operators to specify variations of existing strategies and explore new defense and attack capabilities as they become available. Third, we show how Perry can shed light on the efficacy of various deception defenses and their cost-benefit tradeoffs.
@@ VII-A Reduction in implementation effort
[…]
@@
LLM-based code generation: Given recent advances in LLM-based code generation, a natural question is, if and how, the aforementioned effort can be further reduced using LLMs. As a first-step, we evaluate LLMs ability at generating three deception defenses. To this end, we consider two settings: (1) prompting an LLM to produce code using current tools, low-level APIs, commands, and (2) providing the LLM with Perry APIs and then prompting it to produce code using this API (Fig. 8). For both settings, we use Sonnet 3.7 Thinking as the LLM. The prompts we use are in Appendix  F.
We execute the code to detect errors and also manually review it to validate the implementation. For all three deception strategies, LLMs are able to correctly translate the strategy with Perry’s abstractions. However, the code produced in low-level APIs is incorrect and do not execute without errors. We validated the Sonnet 3.7-generated Perry code against the DFS attacker in the Equifax Large environment and find that the behavior is comparable to the manually written defense code.
@@ VII-B Evaluating Perry’s extensibility
Defender capabilities: As an illustrative case study, we consider the case where an operator wants to explore the benefits of fine-grained host telemetry tools [13, 28] to identify when attackers interact with decoy resources. Specifically, we consider SysFlow [28] to collect system calls using eBPF [13]. This required us to 1) create an Ansible playbook to install SysFlow on each host (26 LOC); 2) add a Python module to configure SysFlow to send the data to the Elasticsearch database (28 LOC); 3) modify environment files to install SysFlow on all hosts (6 LOC); and 4) add an observation module that uses the system calls to detect when attackers use decoy credentials or interact with decoy hosts (42 LOC).
@@
New attacker capabilities: Attackers expressed using the Perry APIs can be extended similar to the defender. Furthermore, Perry can also to support attacker code that does not use the Perry API. For instance, semi- or autonomous LLM-based attackers are starting to show early promise in CTF style challenges [63]. It is easy to plug in such emerging attacker capabilities as well. The effort to support a new autonomous attacker in Perry was quite minimal; we just need to create a new attack strategy that prompts the LLM for bash commands to execute and respond with the results (52 LOC).
@@ VII-C What-If scenarios and interesting findings
Illustrative findings: In what follows, we highlight some interesting findings. We do not intend these as conclusive statements about the state of deception or universal facts across all future environments. We use these as illustrative examples of the types of what-if analysis tradeoffs operators can uncover. For each result, we run the experiment 5 times and report the averages across the runs. In the interest of brevity, for some results we only focus on the environments inspired by the Equifax and Colonial Pipeline incidents.
@@
In the Equifax-inspired environment  we consider: (1) binary success metric if the attacker was able to exfiltrate all data and (b) the time to exfiltrate all data. In the Colonial Pipeline-inspired environment, similarly we use a binary metric if the attacker was able to infect all critical actuators (Fig. 11) and the time the attacker took to infect all actuators (Fig. 9). [^6] For both time metrics, if the attacker failed, we mark the time as the maximum for the environment.
@@
We find static deception strategies can delay but not fundamentally limit the binary success (i.e., infect or exfiltrate critical assets) of many attackers we evaluated. In contrast, stateful deception that reacts to changes in system state are able thwart many types of attacks we considered (Fig. 11).
@@
[…]
@@
[…]
@@
Table VII summarizes running our Sonnet 3.7 Thinking LLM-based attacker for various deception scenarios across 10 trials. (Note that this is different from the use of LLM based code generation for defense strategies in Sec. VII-A). The LLM-based attacker successfully exfiltrated data in 2 out of 10 trials with no deception as shown in Table VII. However, against both deception strategies, it failed to exfiltrate data in any trial. We found that $\approx 90\%$ of commands from the LLM were indeed “wasted” on deceptive resources. While these are the early days in autonomous LLM-based attackers [11, 57], this early result suggests deception may be a promising defense against future autonomous LLM-based attackers.
@@ VIII Discussion and Limitations
Perry could also potentially be used to generate and collect realistic synthetic data to train anomaly detection tools. For example, Perry could run a wide variety of attackers in an environment to train intrusion-detection systems.
@@
Furthermore, a challenge with all emulation tools is the high resource requirements of emulating large networks. For instance, we estimate that emulating a 10,000-host enterprise organization for the same number of experiments in Sec. VII-A would cost \$50,594 on Google Cloud. [^7] We assume each virtual machine has the same requirements as the experiments in Sec. VII-A: 1 CPU, 1GB RAM, and 5GB of disk space. We assume each experiment takes on average 1 hour. To help reduce the cost of evaluating deception on large networks, we have preliminarily evaluated whether experimental results obtained on smaller, simplified environments can carry over to larger, more realistic environments; our experiments suggest that results obtained on smaller environments can carry over to larger environments. As future work, we plan to more thoroughly investigate how to reliably approximate aspects of the environment such as services, vulnerabilities, and human interactions.
@@ X Conclusions
Our work was inspired by the significant obstacles we faced in emulating attackers, deception approaches, and environments with existing tools. Our experiences and challenges helped us identify common design patterns to inform Perry’s programming model for implementing deception what-if experiments. We introduce Perry to demonstrate the importance of what-if experiments for cyber deception. Our work can serve as the basis to provide operators with the tools to explore and optimize deception strategies in their deployment.
Acknowledgments This work was sponsored in part by NSF under award CNS2106214. This work was also supported in part by Microsoft corporation and through the Carnegie Mellon Cylab Future of Enterprise Security initiative. Some of the initial design was also sponsored in part by the Combat Capabilities Development Command Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-13-2-0045 (ARL Cyber Security CRA). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Combat Capabilities Development Command Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes not withstanding any copyright notation here on.